# Assessing the Robustness of Retrieval-Augmented Generation Systems in K-12 Educational Question Answering with Knowledge Discrepancies

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) systems have demonstrated remarkable potential as question answering systems in the K-12 Education domain, where knowledge is typically queried within the restricted scope of authoritative textbooks. However, the discrepancy between textbooks and the parametric knowledge in Large Language Models (LLMs) could undermine the effectiveness of RAG systems. To systematically investigate the robustness of RAG systems under such knowledge discrepancies, we present EduKDQA, a question answering dataset that simulates knowledge discrepancies in real applications by applying hypothetical knowledge updates in answers and source documents. EduKDQA includes 3,005 questions covering five subjects, under a comprehensive question typology from the perspective of context utilization and knowledge integration. We conducted extensive experiments on retrieval and question answering performance. We find that most RAG systems suffer from a substantial performance drop in question answering with knowledge discrepancies, while questions that require integration of contextual knowledge and parametric knowledge pose a challenge to LLMs. All resources will be released to foster further research.
@@
Assessing the Robustness of Retrieval-Augmented Generation Systems in K-12 Educational Question Answering with Knowledge Discrepancies
@@ 1 Introduction
In K-12 educational QA, one of the primary concerns is ensuring that the knowledge conveyed in the answer is consistent with the officially designated textbooks Extance (2023). However, there are notable discrepancies between the knowledge in textbooks and the internal knowledge of LLMs, due to the evolving nature of facts Arbesman (2012), updates in pedagogical approaches Provenzo et al. (2011), as well as regional and cultural variations in content Patel (2015). It remains unclear whether RAG systems can robustly incorporate knowledge from authoritative sources and generate consistent answers under such knowledge discrepancies (scenario illustrated in Figure 1).
To fill this gap, we aim to systematically assess the robustness of RAG systems in performing question answering in K-12 education when encountering knowledge discrepancies. We present EduKDQA (Educational Knowledge Discrepancy Question Answering), a new dataset containing 3,005 multiple-choice questions covering the subjects of Physics, Chemistry, Biology, Geography, and History from the middle-school curriculum. To simulate the knowledge discrepancy between LLMs and textbooks, we conduct a hypothetical knowledge update, in which we modify the original factual knowledge from the textbooks into plausible alternatives while maintaining coherent and consistent context. Moreover, we tailored a comprehensive question typology to stress-test the context utilization and knowledge integration abilities of LLMs under such scenarios.
We conducted extensive experiments with various retrieval methods and LLMs. We find that most RAG systems still suffer from a considerable performance drop when facing knowledge discrepancies. Notably, while most LLMs can incorporate distant contextual facts well, they struggle to integrate their parametric knowledge with contextual knowledge effectively. In terms of retrieval, traditional lexical-based methods show advantages due to specificity of the academic terms, and their performance may be further enhanced through an ensemble reranking mechanism. To encourage further research, we will make our benchmark dataset and the associated code publicly available.
2 The EduKDQA Dataset
In this section, we introduce the methodology of hypothetical knowledge update and the design of our question typology. The curation pipeline and detailed statistics of the EduKDQA dataset are provided in Appendix B and E, respectively.
@@ 2.2 Question Typology
We identify two potential challenges for LLMs when performing QA with knowledge discrepancy: 1) Context Utilization: Whether LLMs can identify and utilize the corresponding facts from the context; and 2) Knowledge Integration: Whether LLMs can incorporate their own parametric knowledge with the contextual ones in question answering. To this end, we designed our question typology, as illustrated in Table 1, to investigate such abilities. Based on the two basic question types, Simple Direct and Multi-hop Direct, we developed the Multi-hop Distant type to evaluate the context utilization ability for distant facts from the passage, and the Multi-hop Implicit111The word ’implicit’ indicates that the questions indirectly query the updated facts by embedding them within the middle of the multi-hop reasoning chain. type to evaluate the knowledge integration ability that combines their own factual knowledge with retrieved ones. Moreover, the Distant Implicit type poses a greater challenge by combining both features. To ensure our evaluation of knowledge integration ability is independent of knowledge coverage, we restrict the facts requiring LLMs’ own knowledge to be head-knowledge only Sun et al. (2024), which we believe are well covered by all tested LLMs.
@@ 3 Experiments and Analyses
Typically, RAG systems first conduct document retrieval based on given queries, then perform question answering with LLMs based on the retrieved information loaded in the context. In this section, we comprehensively evaluate and analyze the performance of retrieval methods and LLMs on the EduKDQA dataset. For details of all tested methods and models, please refer to Appendix C.
@@ 3.1 Retrieval Performance
Moreover, ensemble methods have demonstrated their effectiveness in various information retrieval tasks Thakur et al. (2021). We implemented a hybrid approach that retrieves the top k222We use k = 6 as the optimal hyperparameter setting. documents with Ada-002, followed by re-ranking with BM25. This ensemble method marginally improved the retrieval performance in both metrics. Conversely, query rewriting333We use Mistral-small-2409 as the LLM query rewriter. Ma et al. (2023) implemented upon BM25 did not yield performance gains. Detailed results across subjects and question types are provided in Appendix F.
@@ 3.2 Question Answering Performance
Experimental results of LLMs in question answering are presented in Table 3. Most LLMs exhibited comparable performance for question types Simple Direct, Multi-hop Direct, and Multi-hop Distant. This suggests that both multi-hop reasoning and distant context utilization do not pose significant challenges for modern LLMs. However, for Multi-hop Implicit questions requiring the integration of contextual and internal knowledge, a substantial performance disparity emerges between smaller open-source LLMs and advanced models. This disparity is further amplified for Distant Implicit questions. The most capable model, o1-preview, attains over 80% accuracy on Implicit questions, whereas Mistral-7b’s performance falls below 40%. These findings indicate that knowledge integration is an emergent capability presenting greater difficulties for LLMs under knowledge discrepancies, particularly when coupled with the need for distant context utilization.
@@ 3.3 Overall Performance
How do knowledge discrepancies affect the performance of RAG systems in educational application? To answer this practical question, we selected two representative LLMs: Llama3-8b from open-source models and GPT-4o from commercial models, and combined them with two high-performing retrieval methods: Ada-002 and hybrid rerank. The resulting RAG systems were tested on questions before and after hypothetical knowledge updates, with results presented in Table 4. We observed an accuracy drop of 22-27%, indicating a substantial performance degradation in modern RAG systems when faced with knowledge discrepancies.
@@ 4 Related Work
Prior to the emergence of RAG systems utilizing LLMs, various educational QA systems were developed to provide pedagogically appropriate responses to student inquiries Abdi et al. (2018); Agarwal et al. (2019). While recent literature explores LLM applications in QA and learning assistance roles Nye et al. (2023); Kuo et al. (2023); Wang et al. (2024).
@@
Mitigating knowledge discrepancies or conflicts in applications is a fundamental challenge in LLM research Xu et al. (2024). Researchers have proposed tuning-based Li et al. (2022) and prompting-based Zhou et al. (2023) methods to enhance LLMs’ robustness under such conflicts. However, these conflicts in educational applications and RAG systems remain relatively underexplored.
@@ 5 Conclusion
This paper systematically evaluates the robustness of retrieval-augmented generation systems in K-12 educational question answering under knowledge discrepancies using a comprehensive dataset. Experimental findings reveal substantial performance degradation in RAG systems when faced with knowledge discrepancies, which is primarily attributed to deficiencies in incorporating contextual and parametric knowledge in question answering—an emergent and challenging ability for contemporary large language models.
@@
First, EduKDQA employs the approach of hypothetical knowledge updates to effectively simulate real-world knowledge discrepancies for two primary reasons: (1) Real-world knowledge conflicts are often sparse, noisy, and difficult to systematically collect or organize into a cohesive dataset suitable for large-scale evaluation. Hypothetical updates provide a scalable and controlled alternative, allowing us to bypass these limitations. (2) By leveraging a systematic annotation and curation pipeline, we can generate questions with diverse and well-defined reasoning patterns that align with our typology, enabling more robust evaluation of complex question-answering tasks. For future research centered on real-world knowledge discrepancies, we recommend an alternative methodology that incorporates temporal attributes Chen et al. (2021); Zhang and Choi (2024). This approach focuses on identifying outdated facts and capturing time-sensitive data (e.g., economic trends, annual events, or societal changes) to construct datasets that reflect real-world knowledge updates. While promising, this method is constrained by the limited overlap between time-sensitive data—often numerical or attribute-specific—and the broader contextual needs of educational question-answering tasks, which may reduce the comprehensiveness of the resulting datasets.
@@
In constructing the EduKDQA dataset, we collected text from open-access textbooks. Detailed sources and licenses are provided in Appendix A. The human curation and verification in our annotation pipeline were carried out by a group of postgraduate students with extensive experience in NLP research. We ensured that the updated knowledge is free from harmful or toxic content. It is important to note that our dataset is designed solely to evaluate the robustness of Retrieval-Augmented Generation systems under scenarios with knowledge discrepancies and is not suitable for assessing the factual accuracy of QA systems.
@@ Appendix A Source Textbooks
The documents in the EduKDQA dataset is organized based on the following public textbooks:
@@
Physics by Openstax (CC-BY-4.0444https://creativecommons.org/licenses/by/4.0/deed.en) https://openstax.org/details/books/physics
@@
World History by OER Commons (CC-BY-NC-4.0555https://creativecommons.org/licenses/by-nc/4.0/deed.en) https://oercommons.org/courses/world-history-2
@@
World Regional Geography by Sailor Academy (CC-BY-3.0666https://creativecommons.org/licenses/by/3.0/)
@@
Meanwhile, the detailed topics included in each subjects are illustrated in Figure 2.
@@ Appendix B Curation Pipeline
The curation pipeline of our dataset is illustrated in Figure 3. We first perform triplet extraction on the textbook documents and generate a document-level knowledge graph (KG). Next, we perform sub-graph matching based on fixed reasoning patterns to sample candidate queries, and selectively transform them into natural language questions. Then, hypothetical knowledge update is executed and verified to guarantee consistency between the updated answer and the document. Context-focused question types, including Simple Direct, Multi-hop Direct, and Multi-hop Distant, are acquired through this process. To generate the other two Implicit question types that require knowledge integration, we perform extra QA augmentation followed by an expert verification process. Our data curation process is performed by a mixture of human annotators and LLMs. For LLM annotation, we adopted Claude-3.5-Sonnet for its outstanding instruction-following ability. The total API cost for data annotation is approximately 300 USD.
@@ Appendix C Model Details
In the retrieval stage of our experiments, we employ a diverse range of retrieval methods. For traditional lexical retrieval, we include TF-IDF Spärck Jones (1972), which vectorizes documents and queries based on term frequency and inverse document frequency, and BM25 Robertson et al. (1994), which enhances TF-IDF with document length normalization and probabilistic term weighting. We also incorporate SPLADE Formal et al. (2021), a method that bridges lexical and dense retrieval paradigms. For dense retrieval, we evaluate several methods that encode questions and documents into the same vector space: Contriever Izacard et al. (2022), an unsupervised text encoder; its fine-tuned variant Contriever-msmarco; and two closed-source embedding models, Mistral-embed Mistral (2023a) and Ada-002 OpenAI (2022). Additionally, we explore a pre-retrieval method, Query Rewrite Ma et al. (2023), which reformulates queries to improve retrieval performance. Finally, we implement a Hybrid Rerank approach that combines BM25 and Ada-002, leveraging the strengths of both lexical and dense retrieval methods.
@@ Appendix E Dataset Statistics
The statistics for the documents and questions in the EduKDQA dataset are provided in Table 6 and Table 7, respectively.

### v3

# KnowShiftQA: How Robust are RAG Systems when Textbook Knowledge Shifts in K-12 Education?

@@ Abstract
Retrieval-Augmented Generation (RAG) systems show remarkable potential as question answering tools in the K-12 Education domain, where knowledge is typically queried within the restricted scope of authoritative textbooks. However, discrepancies between these textbooks and the parametric knowledge inherent in Large Language Models (LLMs) can undermine the effectiveness of RAG systems. To systematically investigate RAG system robustness against such knowledge discrepancies, we introduce KnowShiftQA111https://github.com/HKUST-KnowComp/KnowShiftQA. This novel question answering dataset simulates these discrepancies by applying deliberate hypothetical knowledge updates to both answers and source documents, reflecting how textbook knowledge can shift. KnowShiftQA comprises 3,005 questions across five subjects, designed with a comprehensive question typology focusing on context utilization and knowledge integration. Our extensive experiments on retrieval and question answering performance reveal that most RAG systems suffer a substantial performance drop when faced with these knowledge discrepancies. Furthermore, questions requiring the integration of contextual (textbook) knowledge with parametric (LLM) knowledge pose a significant challenge to current LLMs.
@@
KnowShiftQA: How Robust are RAG Systems when Textbook Knowledge Shifts in K-12 Education?
@@ 1 Introduction
In K-12 educational QA, one of the primary concerns is ensuring that the knowledge conveyed in the answer is consistent with the officially designated textbooks Extance (2023). However, notable knowledge shifts (i.e. factuality discrepancies) may exist between the knowledge in textbooks and the internal knowledge of LLMs, due to the evolving nature of facts Arbesman (2012), updates in pedagogical approaches Provenzo et al. (2011), as well as regional and cultural variations in content Patel (2015). It remains unclear whether RAG systems can robustly incorporate knowledge from authoritative sources and generate consistent answers under such knowledge discrepancies (scenario illustrated in Figure 1).
To fill this gap, we aim to systematically assess the robustness of RAG systems in performing question answering in K-12 education when encountering knowledge shifts. We present KnowShiftQA, a new dataset containing 3,005 multiple-choice questions covering the subjects of Physics, Chemistry, Biology, Geography, and History from the middle-school curriculum. To simulate the factuality discrepancy between LLMs and textbooks, we conduct a hypothetical knowledge update, in which we modify the original factual knowledge from the textbooks into plausible alternatives while maintaining coherent and consistent context. Moreover, we tailored a comprehensive question typology to stress-test the context utilization and knowledge integration abilities of LLMs under such scenarios.
We conducted extensive experiments with various retrieval methods and LLMs. Our findings indicate that most RAG systems exhibit a considerable performance degradation under knowledge shifts. Notably, while many LLMs can effectively incorporate distant contextual facts, they struggle to seamlessly integrate their parametric knowledge with this contextual information. In terms of retrieval, traditional lexical-based methods demonstrate advantages due to the specificity of academic terms; their performance can be further enhanced through an ensemble re-ranking mechanism. Supervised retrieval methods (e.g., Contriever), when fine-tuned directly on our dataset, yield superior performance. Overall, our results highlight the fragility of current RAG systems under knowledge shifts and the challenges LLMs face in effectively combining contextual and parametric knowledge for complex question answering.
2 The KnowShiftQA Dataset
In this section, we introduce the methodology of hypothetical knowledge update and the design of our question typology. The curation pipeline and detailed statistics of the KnowShiftQA dataset are provided in Appendix B and E, respectively.
@@ 2.2 Question Typology
We identify two potential challenges for LLMs when performing QA under knowledge shifts: 1) Context Utilization: Whether LLMs can identify and utilize the corresponding facts from the context; and 2) Knowledge Integration: Whether LLMs can incorporate their own parametric knowledge with contextual facts in question answering. To this end, we designed our question typology, as illustrated in Table 1, to investigate such abilities. The two basic question types, Simple Direct and Multi-hop Direct, aim to evaluate LLMs on simple factual recall and multi-hop reasoning. Based on these two types, we developed the Multi-hop Distant type to evaluate the context utilization ability for distant facts from the passage, and the Multi-hop Implicit222The word ’implicit’ indicates that the questions indirectly query the updated facts by embedding them within the middle of the multi-hop reasoning chain. type to evaluate the knowledge integration ability that combines their own factual knowledge with retrieved facts. Moreover, the Distant Implicit type poses a greater challenge by combining both features. To ensure our evaluation of knowledge integration ability is independent of knowledge coverage, we restrict the facts requiring LLMs’ own knowledge to be head-knowledge only Sun et al. (2024).
@@ 3 Experiments and Analyses
Typically, RAG systems first conduct document retrieval based on given queries, then perform question answering with LLMs based on the retrieved information loaded in the context. In this section, we comprehensively evaluate and analyze the performance of retrieval methods and LLMs on the KnowShiftQA dataset. For details of all tested methods and models, please refer to Appendix C.
@@ 3.1 Retrieval Performance
Moreover, ensemble methods have demonstrated their effectiveness in various information retrieval tasks Thakur et al. (2021). We implemented a hybrid approach that retrieves the top k333We use k = 6 as the optimal hyperparameter setting. documents with Ada-002, followed by re-ranking with BM25. This ensemble method marginally improved the retrieval performance in both metrics. Conversely, query rewriting444We use Mistral-small-2409 as the LLM query rewriter. Ma et al. (2023) implemented upon BM25 did not yield performance gains. Detailed results across subjects and question types are provided in Appendix F.
@@ 3.2 Question Answering Performance
Experimental results of LLMs in question answering are presented in Table 3. Most LLMs exhibited comparable performance for question types Simple Direct, Multi-hop Direct, and Multi-hop Distant. This suggests that both multi-hop reasoning and distant context utilization do not pose significant challenges for modern LLMs. However, for Multi-hop Implicit questions, which requires the integration of contextual and internal knowledge, a substantial performance disparity emerges between smaller open-source LLMs and advanced models. This disparity is further amplified for Distant Implicit questions, where the reasoning chains are getting more complexed. The most capable model we tested, o1-preview, attains over 80% accuracy on Implicit questions, whereas Mistral-7b’s performance falls below 40%. These findings indicate that context-memory knowledge integration is an emergent capability presenting greater difficulties for LLMs under knowledge discrepancies, particularly when coupled with the need for distant context utilization.
@@ 3.3 Overall Performance
How do knowledge shifts affect the performance of RAG systems in educational question-answering applications? To answer this question, we selected two representative LLMs: Llama3-8b from open-source models and GPT-4o from proprietary models. These were combined with two high-performing retrieval methods: Ada-002 and hybrid rerank. The resulting RAG systems were tested on questions before and after hypothetical knowledge updates, with the results presented in Table 4. We observed a significant accuracy drop of 22–27% after these knowledge shifts, indicating substantial performance degradation in modern RAG systems when faced with such scenarios.
@@ 4 Related Work
Following the categorization by Gao et al. (2024), the RAG methods employed in our experiment fall into the categories of Naive RAG (lexical, dense) and Advanced RAG (rerank, rewrite). Recently, Modular RAG has emerged to enhance the adaptability and versatility of RAG systems Shao et al. (2023); Asai et al. (2023). Furthermore, recent advances in vector databases Han et al. (2023) and neural graph databases Bai et al. (2025b) highlight the potential for adaptive and generalizable RAG search with structured data Deng et al. (2024); Bai et al. (2025a).
@@
Prior to the emergence of RAG systems utilizing LLMs, various educational QA systems were developed to provide pedagogically appropriate responses to student inquiries Abdi et al. (2018); Agarwal et al. (2019). While recent literature explores LLM applications in QA and learning assistance roles Nye et al. (2023); Kuo et al. (2023); Wang et al. (2024), it is also crucial to leverage logical inference abilities, such as abductive reasoning Bai et al. (2024); Yim et al. (2024); Zheng et al. (2025b) and inductive reasoning Qiu et al. (2024); Li et al. (2025); Zheng et al. (2025a), for understanding students’ learning outcomes in response to their queries.
@@
Mitigating knowledge discrepancies or conflicts within LLM applications is a fundamental challenge in LLM research Xu et al. (2024). Researchers have proposed tuning-based Li et al. (2022) and prompting-based Zhou et al. (2023) methods to enhance LLMs’ robustness to such conflicts. However, the implications of these conflicts in educational applications and RAG systems remain relatively underexplored. Moreover, it is important to consider the trade-off between factuality robustness Yu et al. (2024); Zong et al. (2024); Zheng et al. (2025c) and flexibility in instruction following when such discrepancies are present.
@@ 5 Conclusion
This paper systematically evaluates the robustness of RAG systems in K-12 educational question answering under knowledge discrepancies using a comprehensive dataset KnowShiftQA. Experimental findings reveal substantial performance degradation in RAG systems when faced with knowledge discrepancies, which is primarily attributed to deficiencies in incorporating contextual and parametric knowledge in question answering—an emergent and challenging ability for modern large language models.
@@
First, KnowShiftQA employs the approach of hypothetical knowledge updates to effectively simulate real-world knowledge discrepancies for two primary reasons: (1) Real-world knowledge conflicts are often sparse, noisy, and difficult to systematically collect or organize into a cohesive dataset suitable for large-scale evaluation. Hypothetical updates provide a scalable and controlled alternative, allowing us to bypass these limitations. (2) By leveraging a systematic annotation and curation pipeline, we can generate questions with diverse and well-defined reasoning patterns that align with our typology, enabling more robust evaluation of complex question-answering tasks. For future research centered on real-world knowledge discrepancies, we recommend an alternative methodology that incorporates temporal attributes Chen et al. (2021); Zhang and Choi (2024). This approach focuses on identifying outdated facts and capturing time-sensitive data (e.g., economic trends, annual events, or societal changes) to construct datasets that reflect real-world knowledge updates. While promising, this method is constrained by the limited overlap between time-sensitive data—often numerical or attribute-specific—and the broader contextual needs of educational question-answering tasks, which may reduce the comprehensiveness of the resulting datasets.
@@
In constructing the KnowShiftQA dataset, we collected text from open-access textbooks. Detailed sources and licenses are provided in Appendix A. The human curation and verification in our annotation pipeline were carried out by a group of postgraduate students with extensive experience in NLP research. We ensured that the updated knowledge is free from harmful or toxic content. It is important to note that our dataset is designed solely to evaluate the robustness of Retrieval-Augmented Generation systems under scenarios with knowledge discrepancies and is not suitable for assessing the factual accuracy of QA systems.
Acknowledgments
The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong SAR, China.
@@ Appendix A Source Textbooks
The documents in the KnowShiftQA dataset is organized based on the following public textbooks (the detailed topics included in each subjects are illustrated in Figure 2.):
@@
Physics by Openstax (CC-BY-4.0555https://creativecommons.org/licenses/by/4.0/deed.en) https://openstax.org/details/books/physics
@@
World History by OER Commons (CC-BY-NC-4.0666https://creativecommons.org/licenses/by-nc/4.0/deed.en) https://oercommons.org/courses/world-history-2
@@
World Regional Geography by Sailor Academy (CC-BY-3.0777https://creativecommons.org/licenses/by/3.0/)
@@ Appendix B Curation Pipeline
The curation pipeline of our dataset is illustrated in Figure 3. We first perform triplet extraction on the textbook documents and generate a document-level knowledge graph (KG). Next, we perform sub-graph matching based on fixed reasoning patterns to sample candidate queries, and selectively transform them into natural language questions. Then, hypothetical knowledge update is executed and verified to guarantee consistency between the updated answer and the document.
Context-focused question types, including Simple Direct, Multi-hop Direct, and Multi-hop Distant, are acquired through this process. For Multi-hop Distant questions, we leverage distant facts, defined as connected triplet pairs that are separated in the document’s sequential ordering. These questions are only assigned to documents containing no fewer than 200 words. To generate the other two Implicit question types that require knowledge integration, we perform extra QA augmentation followed by an expert verification process. The design of our question typology is motivated by studies in complex logical reasoning Bai et al. (2023); Zheng et al. (2024, 2025d).
Our data curation process is performed through an integrated framework involving both human annotators and LLMs. For LLM annotation, we adopted Claude-3.5-Sonnet for its outstanding instruction-following ability. Following manual verification, 90.5% of these queries were retained or underwent minor refinements to become high-quality questions, yielding an overall success rate of 86.4%. The total API cost for data annotation is approximately 300 USD.
@@ Appendix C Model Details
In the retrieval stage of our experiments, we employ a diverse range of retrieval methods. For traditional lexical retrieval, we include TF-IDF Spärck Jones (1972), which vectorizes documents and queries based on term frequency and inverse document frequency, and BM25 Robertson et al. (1994), which enhances TF-IDF with document length normalization and probabilistic term weighting. We also incorporate SPLADE Formal et al. (2021), a method that bridges lexical and dense retrieval paradigms. For dense retrieval, we evaluate several methods that encode questions and documents into the same vector space: Contriever Izacard et al. (2022), an unsupervised text encoder; its fine-tuned variant Contriever-msmarco, which we further enhanced by applying contrastive learning Izacard et al. (2022) to fine-tune both models on the KnowShiftQA dataset for improved retrieval capability; and two closed-source embedding models, Mistral-embed Mistral (2023a) and Ada-002 OpenAI (2022). Additionally, we explore a pre-retrieval method, Query Rewrite Ma et al. (2023), which reformulates queries to improve retrieval performance. Finally, we implement a Hybrid Rerank approach that combines BM25 and Ada-002, leveraging the strengths of both lexical and dense retrieval methods.
@@ Appendix E Dataset Statistics
The statistics for the documents and questions in the KnowShiftQA dataset are provided in Table 6 and Table 7, respectively.